• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name V. Sulkosky

    Log entry time 17:04:05 on October 15, 2009

    Entry number 295086

    keyword=Beam Update (Not Good!)

    I went over to MCC and got an update on the beam. Unfortunately just as
    a Mott measurement was underway, there was a network problem that
    brought several of the injector IOC's down. In the process of fixing this,
    a VME power supply had to be replaced which was sparking.

    Even worse, the injector RF phases got hosed. Basically the crew has to
    start over with the injector tune. Mott measured is now postponed for today.

    The relevant ELOG's are attached as figures.

    A copy of this log entry has been emailed to: paschke,camsonne

    ELOG # 1492578

    ELOG # 1492592